OpenStudy (compassionate):

Lets assume f(x) = x^2 -3x-7 Lets say that f(-5) is what we're solving for. It would be f(-5) = -5^2 - 3(-5) - 7, yes?

OpenStudy (camerondoherty):

That's what I would say...

OpenStudy (bibby):

25--15-7=25+15-7=40-7=33 yeah

OpenStudy (compassionate):

So I wouldn't just replace the x, I would the x^2, too.

OpenStudy (dan815):

f(-5) = (-5)^2 - 3(-5) - 7

OpenStudy (bibby):

You're not replacing the x^2. You're replacing the x. x^2 = (x)*(x) -5^2 = (-5)*(-5)=25

OpenStudy (dan815):